OpenStudy (anonymous):

Is (9, -1) a solution of x + 5y = 4?

OpenStudy (anonymous):

Hi DeviousDill, since (9, -1) are in the form of (x, y) what you can try is substitution to see if the answer is equal.

OpenStudy (anonymous):

pluging (9,-1) into x + 5y = 4 yeilds 9 + 5(-1) = 4 does this help?

OpenStudy (anonymous):

Okay, I kind of forgot how to find solutions. I just "plug in" (9) for x and (-1) for y and see if the equation equals out to 4? So...the first number is always x and the second number is always y?

OpenStudy (anonymous):

It would turn out to be, 9 + 5(-1) = 4. So, 9 - 5 = 4. so 4 = 4 ? It turns out equal.